OpenStudy (anonymous):

A rectangular prism is truncated removing 10% of its height. The original height of the prism was 100 meters and it has a length of 4 meters and a width of 20 meters. What is the volume of the truncated prism?

OpenStudy (anonymous):

7,200 meters cubed

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

We must first find the length of the new prism. The new length of our prism can be defined as the length of our prism - the length of our prism multiplied by 1/10 or .10 \[100-(.10*100)=90\] 90 is our new length. We then use our intuition to find the volume of the prism. To calculate the volume we take the bottom and find the area of the square. This will give us an area in meters squared. In order to find the area of the bottom we multiply our length by our width. \[4*20=80\] We now multiply our height by the base to give us meters cubed which is a unit volume. \[80*90=7,200\]